OpenStudy (anonymous):

Two triangles can be formed with the given information. Use the Law of Sines to solve the triangles. A = 55°, a = 12, b = 14

OpenStudy (anonymous):

\[\frac{ \sin A }{ 12 } = \frac{ sinB }{ b}\]

OpenStudy (anonymous):

Plug in all the values and you should be able to solve for B.

OpenStudy (solomonzelman):

after plugging your values in, cross multiply.

OpenStudy (anonymous):

fak you

OpenStudy (anonymous):

B = 38.8°, C = 93.2°, c ≈ 25.8

OpenStudy (anonymous):

FAK

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \bf \frac{Sin(A)}{~a}= \frac{Sin(B)}{b} }\) \(\Huge\color{blue}{ \bf \frac{Sin(55)}{12}= \frac{Sin(B)}{14} }\) solve for B....